OpenStudy (anonymous):

find A4, the fourth term of geometric sequence whose first term is A1 =6720 and common ratio is r=-1/4

OpenStudy (anonymous):

A4= ?

OpenStudy (tkhunny):

a1 = 6720 a2 = 6720*(-1/4) a3 = 6720*(-1/4)^2 etc.